The cutting department of Cassel Company has the following production and cost data for July

1

Transferred out 14,100 units

Beginning work in process

$0

2

Started 5,100 units that are 60% complete as to conversion costs and 100% complete as to materials at July 31.

Materials

74,304

Labor

30,000

Manufacturing overhead

21,300

Materials are entered at the beginning of the process. Conversion costs are incurred uniformly during the process.

Compute unit costs and cost reconciliation schedule
Answer #1
1
Materials Conversion Costs
Total Equivalent units of production 19200 17160
2
Materials Conversion Costs
Unit costs 3.87 2.99
3
Cost reconciliation
Costs accounted for
Transferred out 96726 =14100*6.86
Work in process, July 31
      Materials 19737 =5100*3.87
      Conversion costs 9149 28886 =3060*2.99
Total costs 125612
Workings:
Equivalent Units
Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, July 1 0
Started into production 19200
Total units 19200
Units accounted for
Transferred out 14100 14100 14100
Work in process, July 31 5100 5100 3060
Total units 19200 19200 17160
Materials Conversion Costs Total
Unit costs:
Costs in July 74304 51300 125604
Equivalent units 19200 17160
Unit costs 3.87 2.99 6.86